OpenStudy (anonymous):

3x+6=24-x how can i find this answer

OpenStudy (anonymous):

i have an answer

OpenStudy (anonymous):

add x to each side s you get 4x+6=24 subtract 6 from each side and you get 18 divide by 4 and you get 4.5 :)

OpenStudy (anonymous):

3x+x=24-6 4x=18 x=18/4 x=4.5